1. a For a spring with k = 20.0 N/m and hanging mass 100g, what is...

1.

a For a spring with k = 20.0 N/m and hanging mass 100g, what is the angular frequency, w, of any simple harmonic oscillation process?

b. For the spring system in question 1a, if you displace the hanging mass by 0.20 m and then allow the spring to oscillate, at what speed will the hanging mass pass the equilibrium position during each oscillation? You may assume that the spring is massless and is oscillating in a vacuum on the surface of the earth.
Answer #1

1.

as we know

w^2 = k/m

w^2 = 20 / 0.1

w = 14.142 rad/s

=====

2.

velocity at mean positio6

v = Aw = 0.2* 14.142

v = 2.828 m/s

======

comment before rate in case any doubt, will reply for sure.. goodluck